OpenStudy (yuki):

the base of the solid has a region defined by y=x^2 and y=e^x the plane section perpendicular to the y-axis has a shape of sqrt(x). find the volume of the solid

OpenStudy (anonymous):

3d ?

OpenStudy (anonymous):

i cannot solve for the intervals

OpenStudy (yuki):

come on guys I know you want to show off your skills :) Yes, MathMind, it's in 3D

OpenStudy (anonymous):

oh its 3d ok

OpenStudy (yuki):

we are allowed to use a calculator to find the intersections and integrate

OpenStudy (yuki):

this problem is not solve by hand

OpenStudy (anonymous):

do you know what the intervals are ?

OpenStudy (anonymous):

i think i know how to solve it one second

OpenStudy (yuki):

oops, I guess e^x does not intersect x^2 twice so why don't we use x=2 as our other boundary.\[e^x = x^2 \] x = -.703

OpenStudy (yuki):

so the picture I just provide never happens

OpenStudy (anonymous):

now it looks solvable

OpenStudy (anonymous):

\[\int\limits_{-.703}^{2}(e^x-x^2)dx=[e^x-x^3/3]||(-.703, 2)\]

OpenStudy (anonymous):

\[\int\limits_{-.703}^{2}\sqrt{(e^x-x^2)}dx=\]

OpenStudy (anonymous):

3.192926228

OpenStudy (yuki):

YEEEEEEESSSSSSSSS!!!!!